Factorise. 63a^2 –112b^2 ??????????????????????????

Factorise. 
 63a^2 –112b^2 ??????????????????????????

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

63a^2–112b^2
= 7(9a^2–16b^2)
= 7((3a)^2–(4b)^2)
= 7(3a+4b)(3a-4b)
Using identity: x^2-y^2= (x+y)(x-y)

Thanks